OpenStudy (anonymous):

find all integers of n such that : sqrt(50+sqrt(n)) + sqrt(50-sqrt(n)) is an integer number.

OpenStudy (anonymous):

\[\sqrt{50+\sqrt{n}} + \sqrt{50-\sqrt{n}} = k\]

OpenStudy (rational):

For the given expression to be an integer, it must be the case that both \(50+\sqrt{n}\) and \(50-\sqrt{n}\) are perfect squares : \[50+\sqrt{n}=a^2\\50-\sqrt{n}=b^2\\a\ge b\] add and get \[100=a^2+b^2\] Clearly this can be done in exactly two ways : \((a,b)=(10,0),~(8,6)\) and the corresponding \(n\) values are \(n=2500,~196.\)

OpenStudy (loser66):

I have to log in to give you medal. Admire!!!

OpenStudy (anonymous):

you just miss one solution again @rational , 2016 is works also

OpenStudy (rational):

wow! 2016 is also working, how come i missed that

OpenStudy (anonymous):

yeah there are 3 solutions for n :D n = 196, 2500, 2016

OpenStudy (anonymous):

maybe we can use these formula : \[\sqrt{a+b + 2\sqrt{ab}} = \sqrt{a} + \sqrt{b} \]

OpenStudy (anonymous):

and \[\sqrt{a+b - 2\sqrt{ab}} = \sqrt{a} - \sqrt{b}\]

OpenStudy (anonymous):

with a >= b

OpenStudy (anonymous):